an aeroplane leaves an airport and flies due north at a speed of 1800 km/hr. at the same time, another plane leaves the same airport and flies due west at a speed of 1600 km/hr. how far apart will the two planes be after 2.5 hours?
Answer by mananth(16946)

North bound plane speed = 1800 kmph
in 2.5 hours it will have traveled 1800*2.5 = 4500 km
West bound plane speed = 1600 kmph
in 2.5 hours it will have traveled 1600*2.5 = 4000 km
At any given time they form a right triangle with the starting point
Their distance =d
d^2= 4000^2+4500^2
d^2= 36250000
d=sqrt(36250000)
d= 6020.8 km